5x-2y=20
-7x+3y=-26

Answer by stanbon(75887)

5x-2y = 20
-7x+3y =-26
------------
Modify for elimination::
15x - 6y = 3*20
-14x +6y = 2*-26
---------------------
Add and solve for "x":
x = 8
----
solve for "y":
5x - 2y = 20
40 - 2y = 20
2y = 20
y = 10
-------------
Cheers,
Stan H.
------------